OpenStudy (idealist10):

Find a curve y=y(x) through (1, -1) such that the tangent to the curve at any point (xo, y(xo)) intersects the y axis at yI=(xo)^3.

OpenStudy (idealist10):

Let me show you my work first.

OpenStudy (anonymous):

Wasn't this already asked? http://openstudy.com/users/idealist10#/updates/54e3a0cbe4b0f2455c588421 Now you're given an explicit initial condition.

OpenStudy (idealist10):

\[y-y(x _{0})=m(x-x _{0})\]

OpenStudy (anonymous):

Oh it's y-axis this time...

OpenStudy (idealist10):

(0, (xo)^3) (xo)^3-y(xo)=m(0-xo) (xo)^3-y(xo)=-m(xo) -m=((xo)^3-y(xo))/(xo) m=(y(xo)-(xo)^3)/(xo) at (xo, y(xo)) m=dy/dx=(y-x^3)/x dy/(y-x^3)=dx/x And I have trouble integrating dy/(y-x^3).

OpenStudy (anonymous):

You might have better luck treating this as a linear equation. It doesn't appear to be separable. \[\begin{align*} {x_0}^3-y_0&=y'(x_0)(0-x_0)\\\\ -{x_0}^2&=y'(x_0)-\frac{1}{x_0}y_0\end{align*}\]

OpenStudy (idealist10):

How did you get -(xo)^2=y'(xo)-(1/xo)(yo)?

OpenStudy (anonymous):

\[\begin{align*} {x_0}^3-y_0&=y'(x_0)(0-x_0)\\\\ y_0-{x_0}^3&=y'(x_0)x_0\\\\ -{x_0}^3&=y'(x_0)x_0-y_0\\\\ -{x_0}^2&=y'(x_0)-\frac{1}{x_0}y_0\end{align*}\]

OpenStudy (idealist10):

So from there, what do you do?

OpenStudy (anonymous):

The ODE is linear, so you can go ahead with determining the integrating factor.

OpenStudy (idealist10):

Got it! Thanks!